SPECTROSCOPY IN SUPPORT OF PARITY NONCONSERVATION MEASURMENTS: THE A^{2}Π − X^{2}Σ^{+} (0,0) BAND OF BARIUM MONFLUORIDE

\underline{ANH LE}, SARAH FREY AND TIMOTHY C. STEIMLE , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287.

There is renewed interest in the spectroscopy of heavy metal containing polar radical diatomic molecules because they provide a sensitive venue for detection of parity nonconservation (PNC) either from the determination of the electric dipole moment (EDM) ^{a} of the electron, d_{e}, or detection of the interaction of the anapole moment of the nuclei with the unpaired electron ^{b}. The effects due to d_{e} are nuclear spin independent and studies of both the even and odd nuclear spin isotopologues are relevant. Recently, DeMille et al proposed using an odd isotopologue of barium monofluoride, ^{137}BaF, to measure the nuclear spin-dependent parity non-conservation (NSD-PNC) effect resulting from the interaction of the anapole moment of ^{137}Ba with the unpaired electron of the X^{2}Σ^{+} electronic state. Here we report on the analysis of the field-free spectrum of the A^{2}Π − X^{2}Σ^{+}(0,0) band of ^{137}BaF and an analysis of the ^{137}Ba(I = 3/2) and ^{19}F (I = 1/2) hyperfine interaction in the A^{2}Π state. The hyperfine interaction in the X^{2}Σ^{+} state has been previously characterized from the analysis of the pure rotation spectrum.^{c} The optimal optical transitions for monitoring ^{137}BaF in future PNC measurements will be discussed.

aM. G. Kozlov, A. V. Titov, N. S. Mosyagin, and P. V. Souchko, Phys. Rev. A: At., Mol., and Opt. Phys. 56 R3326, 1997.

bD. DeMille, S. B.Cahn, D.Murphree, D. A. Rahmlow, and M. G.Kozlov Phys. Rev. Letts. 100 023003, 2008.

cCh. Ryzlewicz, H-.U.Schutze-Pahlmann, J. Hoeft and T. Torring, Chem.Phys. 71 389, 1982.